Dired Extra Version 2

For The GNU Emacs 19 Directory Editor

Manual Revision: 2.52 1994/08/09 16:51:31

Lawrence R. Dodd

dodd@roebling.poly.edu

(Based on ‘dired.texi’ by Sebastian Kremer <sk@thp.uni-koeln.de>)

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993, 1994 Free Software Foundation Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation. The file used to create this is called ‘dired-x.texi’, but the original work that was altered to make that file was called ‘dired.texi’ written by Sebastian Kremer.

1 Introduction

This documents the extra features for Dired Mode for GNU Emacs 19. It is derived from version 1.191 of Sebastian Kremer’s ‘dired-x.el’ and is GNU Emacs v19 compatible.

In adopting this ‘dired-x.el’ to GNU Emacs v19 some material that has been incorporated into ‘dired.el’ and ‘dired-aux.el’ of the GNU Emacs 19 distribution has been removed and some material was modified for agreement with the functions in ‘dired.el’ and ‘dired-aux.el’. For example, the code using gmhist history functions was replaced with code using the mini-buffer history now built into GNU Emacs 19. Finally, a few other features have been added and a few more functions have been bound to keys.

Please note that ‘dired-x.el’ and this texinfo file ‘dired-x.texi’ are bundled with GNU Emacs versions 19.23 and later.

1.1 Features

Some features provided by Dired Extra

1. Omitting of uninteresting files from dired listing. See Chapter 3 [Omitting Files in Dired], page 6

2. Local variables for dired directories.

See Chapter 4 [Local Variables], page 9

3. Guessing shell commands in dired buffers. See Chapter 5 [Shell Command Guessing], page 10

4. Running dired command in non-dired buffers. See Chapter 6 [Virtual Dired], page 12

5. Finding a file mentioned in a buffer See Section 8.1 [Find File At Point], page 15

6. Commands using file marking.

See Chapter 7 [Advanced Mark Commands], page 13 ‘dired-x.el’ binds some functions to keys in Dired Mode (See [Key Index], page 22) and also binds C-x C-j and C-x 4 C-j globally to dired-jump (See Chapter 9 [Miscellaneous Commands], page 17). It may also bind C-x C-f and C-x 4 C-f to dired-x-find-file and dired-x-find-file-other-window, respectively (See Section 8.1 [Find File At Point], page 15).

1.2 Technical Details

When loaded this code redefines the following functions of GNU Emacs from ‘dired.el’

  • dired-clean-up-after-deletion
  • dired-find-buffer-nocreate
  • dired-initial-position
  • dired-up-directory

and the following functions from ‘dired-aux.el’

  • dired-add-entry
  • dired-read-shell-command

One drawback is that ‘dired-x.el’ will load ‘dired-aux.el’ as soon as dired is loaded. Thus, the advantage of separating out non-essential dired stuff into ‘dired-aux.el’ and only loading when necessary will be lost when ‘dired-x.el’ is used.

2 Installation

This manual describes the dired features provided by the file ‘dired-x.el’. To take advantage of these features, you must load the file and (optionally) set some variables. In your ‘.emacs’ file in your home directory, or in the system-wide initialization file ‘default.el’ in the ‘site-lisp’ directory, put

(add-hook 'dired-load-hook
          (function (lambda ()
                      (load "dired-x")
                      ;; Set dired-x global variables here.  For example:
                      ;; (setq dired-guess-shell-gnutar "gtar")
                      ;; (setq dired-x-hands-off-my-keys nil)
                      )))
(add-hook 'dired-mode-hook
          (function (lambda ()
                      ;; Set dired-x buffer-local variables here.  For example:
                      ;; (setq dired-omit-files-p t)
                      )))

This will load ‘dired-x.el’ when dired is first invoked (for example, when you first do C-x d).

2.1 Optional Installation Dired Jump

In order to have dired-jump and dired-jump-other-window (See Chapter 9 [Miscellaneous Commands], page 17) work before dired and dired-x have been properly loaded the user should set-up an autoload for these functions. In your ‘.emacs’ file put

;;; Autoload `dired-jump' and `dired-jump-other-window'.
;;; We autoload from FILE dired.el.  This will then load dired-x.el
;;; and hence define `dired-jump' and `dired-jump-other-window'.
(define-key global-map "\C-x\C-j" 'dired-jump)
(define-key global-map "\C-x4\C-j" 'dired-jump-other-window)
(autoload (quote dired-jump) "dired" "\
Jump to dired buffer corresponding to current buffer.
If in a file, dired the current directory and move to file's line.
If in dired already, pop up a level and goto old directory's line.
In case the proper dired file line cannot be found, refresh the dired
buffer and try again." t nil)
(autoload (quote dired-jump-other-window) "dired" "\
Like \\[dired-jump] (dired-jump) but in other window." t nil)

Note that in recent releases of GNU Emacs 19 (i.e., 19.25 or later) the file ‘../lisp/loaddefs.el’ of the Emacs distribution already contains the proper auto-loading for dired-jump so you need only put

(define-key global-map "\C-x\C-j" 'dired-jump)

in your ‘.emacs’ file in order to have C-x C-j work before dired is loaded.

2.2 Optional Installation File At Point

   If you choose to have ‘dired-x.el’ bind dired-x-find-file over find-file (See Sec-

tion 8.1 [Find File At Point], page 15), then you will need to set dired-x-hands-off-my-

keys and make a call to the function dired-x-bind-find-file in the dired-load-hook:
     (add-hook 'dired-load-hook
               (function (lambda ()
                           (load "dired-x")
                           ;; Bind dired-x-find-file.
                           (setq dired-x-hands-off-my-keys nil)
                           ;; Make sure our binding preference is invoked.
                           (dired-x-bind-find-file)
                           )))

Alternatively, you can set the variable before ‘dired-x.el’ is loaded

     (add-hook 'dired-load-hook
               (function (lambda ()
                           ;; Bind dired-x-find-file.
                           (setq dired-x-hands-off-my-keys nil)
                           (load "dired-x")
                           )))

2.3 Special Notes

If ‘dired-x.el’ was not bundled with the version of GNU Emacs installed at your site (i.e., not in the default ‘../lisp’ directory) then you must put the file ‘dired-x.el’ in a directory known to GNU Emacs. Examine the variable load-path for a list of these directories. If you wish to add a new directory on this list of directories use something like this in your ‘.emacs’ file

     ;;; LOAD PATH
     (setq load-path (append
                      load-path ; default at top
                      (list
                       "/the/directory/where/you/put/dired-x")))

If you wish to put the new directory at the head of the list (where it will be found first) then you should use instead

     ;;; LOAD PATH
     (setq load-path (append
                      (list
                       "/the/directory/where/you/put/dired-x")
                      load-path)) ; default at bottom

You must also byte compile the file (for example, hitting B in dired-mode). When byte-compiling ‘dired-x.el’ you may get messages about functions vm-visit-folder,

Man-notify-when-ready, and reporter-submit-bug-report not being defined. These

are warnings and should be ignored. CAUTION: If you are using a version of GNU Emacs earlier than 19.20 than you may have to edit ‘dired.el’. The copy of ‘dired.el’ in GNU Emacs versions earlier than 19.20 incorrectly had the call to run-hooks before the call to provide. In such a case, it is possible that byte-compiling and/or loading dired can cause an infinite loop. To prevent this, make sure the line of code

             (run-hooks 'dired-load-hook)

is the last executable line in the file ‘dired.el’. That is, make sure it comes after the line

             (provide 'dired)

3 Omitting Files in Dired

Omitting a file means removing it from the directory listing. Omitting is useful for keeping Dired buffers free of “uninteresting” files (for instance, auto-save, auxiliary, backup, and revision control files) so that the user can concentrate on the interesting files. Like hidden files, omitted files are never seen by Dired. Omitting differs from hiding in several respects:

• Omitting works on individual files, not on directories; an entire directory cannot be omitted (though each of its files could be).

• Omitting is wholesale; if omitting is turned on for a dired buffer, then all uninteresting files listed in that buffer are omitted. The user does not omit (or unomit) files one at a time.

• Omitting can be automatic; uninteresting file lines in the buffer can be removed before the user ever sees them.

• Marked files are never omitted.

M-o (dired-omit-toggle) Toggle between displaying and omitting “uninteresting” files. With a prefix argument, don’t toggle and just mark the files, but don’t actually omit them.

In order to make Dired Omit work you first need to load ‘dired-x.el’ inside dired-load-hook (See Chapter 2 [Installation], page 3) and then set dired-omit-files-p in some way (See Section 3.1 [Omitting Variables], page 6).

3.1 Omitting Variables

The following variables can be used to customize omitting.

dired-omit-files-p

Default: nil If non-nil, “uninteresting” files are not listed. Uninteresting files are those whose filenames match regexp dired-omit-files, plus those ending with ex-

tensions in dired-omit-extensions. M-o (dired-omit-toggle) toggles its

value, which is buffer-local. Put

     (setq dired-omit-files-p t)

inside your dired-mode-hook to have omitting initially turned on in every Dired buffer (See Chapter 2 [Installation], page 3). You can then use M-o to unomit in that buffer. To enable omitting automatically only in certain directories one can use Dired Local Variables and put

     Local Variables:
     dired-omit-files-p: t
     End:

into a file ‘.dired’ (the default value of dired-local-variables-file) in that directory (See Chapter 4 [Local Variables], page 9).

dired-omit-here-always

This is an interactive function that creates a local variables file exactly like the example above (if it does not already exist) in the file dired-local-variables-file in the current directory and then refreshes the directory listing (See Chapter 4 [Local Variables], page 9).

dired-omit-files
Default: "^#\\|\\.$"

Filenames matching this buffer-local regexp will not be displayed. This only

has effect when dired-omit-files-p is t.

The default value omits the special directories ‘.’ and ‘..’ and autosave files (plus other files ending in “.”) (See Section 3.2 [Omitting Examples], page 7).

dired-omit-extensions

Default: The elements of completion-ignored-extensions (as defined in the file ‘loaddefs.el’ of the GNU Emacs distribution), dired-

latex-unclean-extensions, dired-bibtex-unclean-extensions and
dired-texinfo-unclean-extensions.

If non-nil, a list of extensions (strings) to omit from Dired listings. Its format

is the same as that of completion-ignored-extensions.
dired-omit-localp

Default: 'no-dir

The localp argument dired-omit-expunge passes to dired-get-filename. If

it is 'no-dir, omitting is much faster, but you can only match against the non-directory part of the filename. Set it to nil if you need to match the whole pathname or t to match the pathname relative to the buffer’s top-level directory.

dired-omit-marker-char

Default: C-o Temporary marker used by by Dired to implement omitting. Should never be used as marker by the user or other packages. There is one exception to this rule: by doing

     (setq dired-mark-keys "\C-o")
     ;; i.e., the value of dired-omit-marker-char
     ;; (which is not defined yet)

anywhere in your ‘~/.emacs’, you will bind the C-o key to insert a C-o marker, thus causing these files to be omitted in addition to the usually omitted files. Unfortunately the files you omitted manually this way will show up again after reverting the buffer, unlike the others.

3.2 Examples of Omitting Various File Types

• If you wish to avoid seeing RCS files and the RCS directory, then put

     (setq dired-omit-files
           (concat dired-omit-files "\\|^RCS$\\|,v$"))

in the dired-load-hook (See Chapter 2 [Installation], page 3). This assumes diredomit-localp has its default value of 'no-dir to make the ^-anchored matches work. As a slower alternative, with dired-omit-localp set to nil, you can use / instead of ^in the regexp.

• If you use tib, the bibliography program for use with TEX and LaTEX, you might want

to omit the ‘INDEX’ and the ‘-t.tex’ files, then put

     (setq dired-omit-files
           (concat dired-omit-files "\\|^INDEX$\\|-t\\.tex$"))

in the dired-load-hook (See Chapter 2 [Installation], page 3).

• If you do not wish to see ‘dot’ files (files starting with a ‘.’), then put

     (setq dired-omit-files
           (concat dired-omit-files "\\|^\\..+$"))

in the dired-load-hook (See Chapter 2 [Installation], page 3).

3.3 Some Technical Details of Omitting

Loading ‘dired-x.el’ will install Dired Omit by putting dired-omit-expunge on

your dired-after-readin-hook, and will call dired-extra-startup, which in turn calls
dired-omit-startup in your dired-mode-hook.

4 Local Variables for Dired Directories

When Dired visits a directory, it looks for a file whose name is the value of variable diredlocal-variables-file (default: ‘.dired’). If such a file is found, Dired will temporarily insert it into the Dired buffer and run hack-local-variables. For example, if the user puts

     Local Variables:
     dired-actual-switches: "-lat"
     dired-omit-files-p: t
     End:

into a file called ‘.dired’ in a directory then when that directory is viewed it will be

1. sorted by date

2. omitted automatically

You can set dired-local-variables-file to nil to suppress this. The value of diredenable-local-variables controls if and how these local variables are read. This variable exists so that if may override the default value of enable-local-variables. Please see the GNU Emacs Manual to learn more about local variables. See section “Local Variables in Files” in The GNU Emacs Manual. The following variables affect Dired Local Variables

dired-local-variables-file
          Default: ".dired"

If non-nil, filename for local variables for Dired. If Dired finds a file with that name in the current directory, it will temporarily insert it into the dired buffer and run ‘hack-local-variables’.

dired-enable-local-variables

Default: t

Controls use of local-variables lists in dired. The value can be t, nil, or something else. A value of t means local-variables lists are obeyed in the dired-local-variables-file; nil means they are ignored; anything else means query. This variable temporarily overrides the value of enable-local-variables when the Dired Local Variables are hacked.

5 Shell Command Guessing

Based upon the name of a filename, Dired tries to guess what shell command you might want to apply to it. For example, if you have point on a file named ‘foo.tar’ and you press !, Dired will guess you want to ‘tar xvf’ it and suggest that as the default shell command.

The default will be mentioned in brackets and you can type M-p to get the default into the minibuffer so that you can edit it, e.g., changing ‘tar xvf’ to ‘tar tvf’. If there are several commands for a given file, e.g., ‘xtex’ and ‘dvips’ for a ‘.dvi’ file, you can type M-p several times to see each of the matching commands.

Dired only tries to guess a command for a single file, never for a list of marked files.

dired-guess-shell-alist-default

Predefined rules for shell commands. Set this to nil to turn guessing off. The elements of dired-guess-shell-alist-user (defined by the user) will override these rules.

dired-guess-shell-alist-user

If non-nil, a user-defined alist of file regexps and their suggested commands. These rules take precedence over the predefined rules in the variable dired-

guess-shell-alist-default (to which they are prepended) when dired-do-
shell-command is run).

Each element of the alist looks like

(regexp command...) where each command can either be a string or a lisp expression that evaluates to a string. If several COMMANDs are given, all will temporarily be pushed on the history. You can set this variable in your ‘~/.emacs’. For example, to add rules for ‘.foo’ and ‘.bar’ file extensions, write

     (setq dired-guess-shell-alist-user
           (list
            (list "\\.foo$" "foo-command");; fixed rule
            ;; possibly more rules...
            (list "\\.bar$";; rule with condition test

'(if condition

"bar-command-1" "bar-command-2"))))

This will override any predefined rules for the same extensions.

dired-guess-shell-gnutar

Default: nil If non-nil, name of the GNU tar executable (e.g., ‘"tar"’ or ‘"gnutar"’). GNU tar’s ‘z’ switch is used for compressed tar files. If you don’t have GNU tar, set this to nil: a pipe using ‘zcat’ is then used.

dired-guess-shell-gzip-quiet

Default: t A non-nil value means that -q is passed to gzip overriding a verbose GNU zip’s ‘GZIP’ environment variable.

dired-guess-shell-znew-switches nil

Default: nil A string of switches passed to GNU zip’s ‘znew’. An example is ‘"-K"’ which will make ‘znew’ keep a .Z file when it is smaller than the .gz file.

dired-shell-command-history nil

History list for commands that read dired-shell commands.

6 Virtual Dired

Using Virtual Dired means putting a buffer with Dired-like contents in Dired mode. The files described by the buffer contents need not actually exist. This is useful if you want to peruse an ‘ls -lR’ output file, for example one you got from an FTP server. You can use all motion commands usually available in Dired. You can also use it to save a Dired buffer in a file and resume it in a later session.

Type M-x dired-virtual to put the current buffer into virtual Dired mode. You will be prompted for the top level directory of this buffer, with a default value guessed from the buffer contents. To convert the virtual to a real Dired buffer again, type g (which calls dired-virtual-revert) in the virtual Dired buffer and answer ‘y’. You don’t have to do this, though: you can relist single subdirectories using l (dired-do-redisplay) on the subdirectory headerline, leaving the buffer in virtual Dired mode all the time.

The function ‘dired-virtual-mode’ is specially designed to turn on virtual Dired mode from the auto-mode-alist. To edit all ‘*.dired’ files automatically in virtual Dired mode, put this into your ‘~/.emacs’:

     (setq auto-mode-alist (cons '("[^/]\\.dired$" . dired-virtual-mode)
                                   auto-mode-alist))

The regexp is a bit more complicated than usual to exclude ".dired" local variable files.

7 Advanced Mark Commands

F (dired-do-find-marked-files) Find all marked files at once displaying simultaneously. If optional NOSELECT is non-nil then just find the files but do not select. If you want to keep the dired buffer displayed, type C-x 2 first. If you want just the marked files displayed and nothing else, type C-x 1 first. The current window is split across all files marked, as evenly as possible. Remaining lines go to the bottom-most window. The number of files that can be displayed this way is restricted by the height of the current window and the

variable window-min-height.
dired-mark-extension

Mark all files with a certain extension for use in later commands. A ‘.’ is not automatically prepended to the string entered. When called from lisp, extension may also be a list of extensions and an optional argument marker-char specifies the marker used.

dired-flag-extension

Flag all files with a certain extension for deletion. A ‘.’ is not automatically prepended to the string entered.

7.1 Advanced Cleaning Functions

dired-clean-patch

Flag dispensable files created by the ‘patch’ program for deletion. See variable

dired-patch-unclean-extensions.
dired-clean-tex

Flag dispensable files created by TEX, LaTEX, and ‘texinfo’ for deletion. See the following variables (See Section 7.2 [Advanced Cleaning Variables], page 13)

  • dired-tex-unclean-extensions
  • dired-texinfo-unclean-extensions
  • dired-latex-unclean-extensions
  • dired-bibtex-unclean-extensions
dired-very-clean-tex

Flag dispensable files created by TEX, LaTEX, ‘texinfo’, and ".dvi" files for deletion.

7.2 Advanced Cleaning Variables

Variables used by the above cleaning commands (and in the default value for variable dired-omit-extensions, See Section 3.1 [Omitting Variables], page 6)

dired-patch-unclean-extensions
Default: '(".rej" ".orig")

List of extensions of dispensable files created by the ‘patch’ program.

dired-tex-unclean-extensions
Default: '(".toc" ".log" ".aux")

List of extensions of dispensable files created by TEX.

dired-texinfo-unclean-extensions
Default: '(".cp" ".cps" ".fn" ".fns" ".ky" ".kys" ".pg" ".pgs" ".tp"
".tps" ".vr" ".vrs")

List of extensions of dispensable files created by ‘texinfo’.

dired-latex-unclean-extensions
Default: '(".idx" ".lof" ".lot" ".glo")

List of extensions of dispensable files created by LaTEX.

dired-bibtex-unclean-extensions
Default: '(".blg" ".bbl")

List of extensions of dispensable files created by BibTEX.

7.3 Special Marking Function

M-( (dired-mark-sexp) Mark files for which predicate returns non-nil. With a prefix argument, unflag those files instead. The predicate is a lisp expression that can refer to the following symbols: inode [integer] the inode of the file (only for ‘ls -i’ output) s [integer] the size of the file for ‘ls -s’ output (usually in blocks or,

with ‘-k’, in KBytes)

mode [string] file permission bits, e.g., ‘"-rw-r--r--"’ nlink [integer] number of links to file uid [string] owner gid [string] group (If the gid is not displayed by ‘ls’, this will still be

set (to the same as uid))

size [integer] file size in bytes time [string] the time that ‘ls’ displays, e.g., ‘"Feb 12 14:17"’ name [string] the name of the file sym [string] if file is a symbolic link, the linked-to name, else ‘""’ For example, use

     (equal 0 size)

to mark all zero length files. To find out all not yet compiled Emacs lisp files in a directory, dired all ‘.el’ files in the lisp directory using the wildcard ‘*.el’. Then use M-( with

     (not (file-exists-p (concat name "c")))

to mark all ‘.el’ files without a corresponding ‘.elc’ file.

8 Multiple Dired Directories and Non-Dired Commands

An Emacs buffer can have but one working directory, stored in the buffer-local variable default-directory. A Dired buffer may have several subdirectories inserted, but still has but one working directory: that of the top level Dired directory in that buffer. For some commands it is appropriate that they use the current Dired directory instead of default-

directory, e.g., find-file and compile.

A general mechanism is provided for special handling of the working directory in special major modes:

default-directory-alist
Default: ((dired-mode . (dired-current-directory)))

Alist of major modes and their opinion on default-directory, as a lisp expression to evaluate. A resulting value of nil is ignored in favor of default-

directory.
default-directory

Function with usage like variable default-directory, but knows about the

special cases in variable default-directory-alist.

8.1 Find File At Point

‘dired-x’ provides a method of visiting or editing a file mentioned in the buffer you are viewing (e.g., a mail buffer, a news article, a README file, etc.) or to test if that file exists. You can then modify this in the minibuffer after snatching the filename.

When installed ‘dired-x’ will substitute dired-x-find-file for find-file (normally

bound to C-x C-f) and dired-x-find-file-other-window for find-file-other-window

(normally bound to C-x 4 C-f).

In order to use this feature, you will need to set dired-x-hands-off-my-keys to nil inside dired-load-hook (See Section 2.2 [Optional Installation File At Point], page 4).

dired-x-find-file

dired-x-find-file behaves exactly like find-file (normally bound to C-x C-f) unless a prefix argument is passed to the function in which case it will use the filename at point as a guess for the file to visit. For example, if the buffer you were reading contained the words

     Available via anonymous ftp in
        /roebling.poly.edu:/pub/lisp/crypt++.el.gz

then you could move your cursor to the line containing the ftp address and type C-u C-x C-f (the C-u is a universal argument). The minibuffer would read

     Find file: /roebling.poly.edu:/pub/lisp/crypt++.el.gz

with the point after the last /. If you hit return emacs will visit the file at that address. This also works with files on your own computer.

dired-x-find-file-other-window
dired-x-find-file-other-window behaves exactly like find-file-other-

window (normally bound to C-x 4 C-f) unless a prefix argument is used. See dired-x-find-file for more information.

dired-x-hands-off-my-keys

If set to t, then it means that ‘dired-x’ should not bind dired-x-find-file over find-file on keyboard. Similarly, it should not bind dired-x-find-

file-other-window over find-file-other-window. If you change this vari-
able after ‘dired-x.el’ is loaded then do M-x dired-x-bind-find-file. The

default value of this variable is t; by default, the binding is not done. See See Section 2.2 [Optional Installation File At Point], page 4.

dired-x-bind-find-file

A function, which can be called interactively or in your ‘~/.emacs’ file, that

uses the value of dired-x-hands-off-my-keys to determine if dired-x-find-
file should be bound over find-file and dired-x-find-file-other-window

bound over find-file-other-window. See See Section 2.2 [Optional Installation File At Point], page 4.

9 Miscellaneous Commands

Miscellaneous features not fitting anywhere else:

dired-find-subdir

Default: nil If non-nil, Dired does not make a new buffer for a directory if it can be found (perhaps as subdirectory) in some existing Dired buffer. If there are several Dired buffers for a directory, the most recently used is chosen. Dired avoids switching to the current buffer, so that if you have a normal and a wildcard buffer for the same directory, C-x d RET will toggle between those two.

M-g (dired-goto-file) Goto file line of a file (or directory).

M-G (dired-goto-subdir) Goto headerline of an inserted directory. This commands reads its argument with completion over the names of the inserted subdirectories.

w  (dired-copy-filename-as-kill) The w command puts the names of the

marked (or next N) files into the kill ring, as if you had killed them with C-w. With a zero prefix argument N=0, use the complete pathname of each file. With a raw (just C-u) prefix argument, use the relative pathname of each marked file. As a special case, if no prefix argument is given and point is on a directory headerline, it gives you the name of that directory, without looking for marked files. The list of names is also stored onto the variable dired-marked-files for use, e.g., in the M-: (eval-expression) command. As this command also displays what was pushed onto the kill ring you can use it to display the list of currently marked files in the echo area (unless you happen to be on a subdirectory headerline). You can then feed the file name to other Emacs commands with C-y. For example, say you want to rename a long filename to a slightly different name. First type w to push the old name onto the kill ring. Then type R to rename it and use C-y inside R’s minibuffer prompt to insert the old name at a convenient place.

T (dired-do-toggle) Toggle marks. That is, currently marked files become unmarked and vice versa. Files marked with other flags (such as ‘D’) are not affected. The special directories ‘.’ and ‘..’ are never toggled.

dired-smart-shell-command

Like function shell-command, but in the current Dired directory. Bound to M-! in Dired buffers.

dired-jump

Bound to C-x C-j. Jump back to dired: If in a file, dired the current directory and move to file’s line. If in Dired already, pop up a level and goto old directory’s line. In case the proper Dired file line cannot be found, refresh the Dired buffer and try again.

dired-jump-other-window

Bound to C-x 4 C-j. Like dired-jump, but to other window. These functions can be autoloaded so they work even though ‘dired-x.el’ has not been loaded yet (See Section 2.1 [Optional Installation Dired Jump], page 3). If the variable dired-bind-jump is nil, dired-jump will not be bound to C-x

C-j and dired-jump-other-window will not be bound to C-x 4 C-j.

dired-vm Bound to V if dired-bind-vm is t. Run VM on this file (assumed to be a UNIX mail folder). If you give this command a prefix argument, it will visit the folder read-only. This only works in VM~5, not VM~4. If the variable dired-vm-read-only-folders is t, dired-vm will visit all folders read-only. If it is neither nil nor t, e.g., the symbol 'if-file-read-only, only files not writable by you are visited read-only. This is the recommended value if you run VM 5. If the variable dired-bind-vm is t, dired-vm will be bound to V. Otherwise,

dired-bind-rmail will be bound.
dired-rmail

Bound to V if dired-bind-vm is nil. Run Rmail on this file (assumed to be mail folder in Rmail/BABYL format).

dired-info

Bound to I. Run Info on this file (assumed to be a file in Info format). If the variable dired-bind-info is nil, dired-info will not be bound to I.

dired-man

Bound to N. Run man on this file (assumed to be a file in nroff format). If the variable dired-bind-man is nil, dired-man will not be bound to N.

dired-do-relative-symlink

Bound to Y. Relative symlink all marked (or next ARG) files into a directory, or make a relative symbolic link to the current file. This creates relative symbolic links like foo -> ../bar/foo not absolute ones like foo -> /ugly/path/that/may/change/any/day/bar/foo

dired-do-relative-symlink-regexp

Bound to %Y. Relative symlink all marked files containing REGEXP to NEW-NAME. See functions ‘dired-do-rename-regexp’ and ‘dired-do-relsymlink’ for more info.

10 Bugs

If you encounter a bug in this package, wish to suggest an enhancement, or want to make a smart remark, then type

     M-x dired-x-submit-report

to set up an outgoing mail buffer, with the proper address to the ‘dired-x.el’ maintainer automatically inserted in the ‘To:’ field. This command also inserts information that the Dired X maintainer can use to recreate your exact setup, making it easier to verify your bug or social maladjustment.

Lawrence R. Dodd <dodd@roebling.poly.edu> Concept Index 20

Concept Index

A

Adding to the kill ring in dired. . . . . . . . . . . . . . . . 17 Autoloading dired-jump and

dired-jump-other-window . . . . . . . . . . . . . . . . 3

B

Binding dired-x-find-file . . . . . . . . . . . . . . . . . . . 4 Bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

D

‘dired-aux.el’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Dot files, how to omit them in Dired . . . . . . . . . . . 8

F

Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Finding a file at point . . . . . . . . . . . . . . . . . . . . . . . . 15

G

GNU zip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 11 Guessing shell commands for files.. . . . . . . . . . . . . 10

H

How to make omitting the default in Dired . . . . . 6

J

Jumping to dired listing containing file.. . . . . . . . 17

L

Lisp expression, marking files with in Dired …. 14 Local Variables for Dired Directories . . . . . . . . . . . 9 ls listings, how to peruse them in Dired. . . . . . . . 12

M

Mark file by lisp expression . . . . . . . . . . . . . . . . . . . 14 Multiple Dired directories . . . . . . . . . . . . . . . . . . . . 15

O

Omitting additional files. . . . . . . . . . . . . . . . . . . . . . . 7 Omitting dot files in Dired . . . . . . . . . . . . . . . . . . . . 8 Omitting Files in Dired . . . . . . . . . . . . . . . . . . . . . . . 6 Omitting RCS files in Dired . . . . . . . . . . . . . . . . . . . 7 Omitting tib files in Dired . . . . . . . . . . . . . . . . . . . . . 8

P

Passing GNU tar its ‘z’ switch. . . . . . . . . . . . . . . . 10 Perusing ls listings . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

R

RCS files, how to omit them in Dired. . . . . . . . . . . 7 Reading mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Redefined functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Relative symbolic links. . . . . . . . . . . . . . . . . . . . . . . 18 Running info.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Running man. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

S

Simultaneous visiting of several files . . . . . . . . . . . 13

T

Tib files, how to omit them in Dired . . . . . . . . . . . 8 Toggling marks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

V

Virtual Dired . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Visiting a file mentioned in a buffer . . . . . . . . . . . 15 Visiting several files at once . . . . . . . . . . . . . . . . . . 13

W

Working directory . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

Function Index 21

Function Index

D

default-directory . . . . . . . . . . . . . . . . . . . . . . . . . . 15 dired-clean-patch . . . . . . . . . . . . . . . . . . . . . . . . . . 13 dired-clean-tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

dired-copy-filename-as-kill . . . . . . . . . . . . . . . 17
dired-do-find-marked-files . . . . . . . . . . . . . . . . 13
dired-do-relative-symlink . . . . . . . . . . . . . . . . . 18
dired-do-relative-symlink-regexp. . . . . . . . . . 18

dired-do-toggle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 dired-flag-extension. . . . . . . . . . . . . . . . . . . . . . . 13 dired-goto-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 dired-goto-subdir . . . . . . . . . . . . . . . . . . . . . . . . . . 17 dired-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 dired-jump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

dired-jump-other-window . . . . . . . . . . . . . . . . . . . 18

dired-man . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 dired-mark-extension. . . . . . . . . . . . . . . . . . . . . . . 13 dired-mark-sexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

                                                dired-omit-here-always . . . . . . . . . . . . . . . . . . . . . 6

dired-omit-toggle . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 dired-rmail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

                                                dired-smart-shell-command . . . . . . . . . . . . . . . . . 17

dired-very-clean-tex. . . . . . . . . . . . . . . . . . . . . . . 13 dired-virtual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 dired-virtual-mode . . . . . . . . . . . . . . . . . . . . . . . . . 12 dired-virtual-revert. . . . . . . . . . . . . . . . . . . . . . . 12 dired-vm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

                                                dired-x-bind-find-file . . . . . . . . . . . . . . . . . . . . 16

dired-x-find-file . . . . . . . . . . . . . . . . . . . . . . . . . . 15

                                                dired-x-find-file-other-window . . . . . . . . . . . . 15

dired-x-submit-report . . . . . . . . . . . . . . . . . . . . . 19

S

shell-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Key Index 22

Key Index

%

%Y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

C

C-x 4 C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-x 4 C-j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 C-x C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-x C-j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

F

F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

G

g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

I

I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

M

M-! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 M-( . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 M-g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 M-G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 M-o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

N

N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

T

T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

V

V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

W

w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Y

Y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

Variable Index 23

Variable Index

A

auto-mode-alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

D

default-directory-alist . . . . . . . . . . . . . . . . . . . 15
dired-bibtex-unclean-extensions . . . . . . . . . . . 14

dired-bind-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 dired-bind-jump . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 dired-bind-man . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 dired-bind-vm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

dired-enable-local-variables . . . . . . . . . . . . . . . 9

dired-find-subdir . . . . . . . . . . . . . . . . . . . . . . . . . . 17

dired-guess-shell-alist-default . . . . . . . . . . . 10
dired-guess-shell-alist-user . . . . . . . . . . . . . . 10
dired-guess-shell-gnutar . . . . . . . . . . . . . . . . . . 10

dired-guess-shell-gzip-quiet . . . . . . . . . . . . . . 10
dired-guess-shell-znew-switches nil . . . . . . . 11
dired-latex-unclean-extensions . . . . . . . . . . . . 14
dired-local-variables-file . . . . . . . . . . . . . . . . . 9

dired-marked-files . . . . . . . . . . . . . . . . . . . . . . . . . 17 dired-omit-extensions. . . . . . . . . . . . . . . . . . . . . . . 7 dired-omit-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 dired-omit-files-p . . . . . . . . . . . . . . . . . . . . . . . . . . 6 dired-omit-localp . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

dired-omit-marker-char . . . . . . . . . . . . . . . . . . . . . 7
dired-patch-unclean-extensions . . . . . . . . . . . . 13
dired-shell-command-history nil . . . . . . . . . . . 11
dired-tex-unclean-extensions . . . . . . . . . . . . . . 14
dired-texinfo-unclean-extensions. . . . . . . . . . 14
dired-vm-read-only-folders . . . . . . . . . . . . . . . . 18
dired-x-hands-off-my-keys . . . . . . . . . . . . . . . . . 16

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Technical Details . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Optional Installation Dired Jump. . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Optional Installation File At Point . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Special Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Omitting Files in Dired . . . . . . . . . . . . . . . . . . . . 6

3.1 Omitting Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Examples of Omitting Various File Types . . . . . . . . . . . . . . . . . 7 3.3 Some Technical Details of Omitting. . . . . . . . . . . . . . . . . . . . . . . 8

4 Local Variables for Dired Directories . . . . . . . . 9 5 Shell Command Guessing . . . . . . . . . . . . . . . . . 10 6 Virtual Dired. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 7 Advanced Mark Commands . . . . . . . . . . . . . . . 13

7.1 Advanced Cleaning Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.2 Advanced Cleaning Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.3 Special Marking Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

8 Multiple Dired Directories and Non-Dired Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

8.1 Find File At Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

9 Miscellaneous Commands . . . . . . . . . . . . . . . . . 17 10 Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23